OpenStudy (anonymous):

the speed of a remote controlled car is 0.2 kilometers per minute. if there is no reason for the car to stop traveling, how many miles will it travel in 2 hours and 15 minutes?

ganeshie8 (ganeshie8):

we use this formula :- distance = speed x time

ganeshie8 (ganeshie8):

and convert time to same units (minutes would be good)

ganeshie8 (ganeshie8):

distance = speed x time = \(\frac{0.2 km}{minute} \times \text{2 hours 15 minutes}\) = \(\frac{0.2 km}{minute} \times \text{2 * 60 minutes + 15 minutes}\) = \(\frac{0.2 km}{minute} \times \text{135 minutes}\)

ganeshie8 (ganeshie8):

multiply and simplify

OpenStudy (anonymous):

i dont get it @ganeshie8 i got 33.75

ganeshie8 (ganeshie8):

how ?

ganeshie8 (ganeshie8):

looks wrong

ganeshie8 (ganeshie8):

0.2 x 135 = 27 km

OpenStudy (anonymous):

wait what am i supposed to do then?

ganeshie8 (ganeshie8):

convert it to miles, that gives :- \(27 km \times \frac{1 mile}{1.609 km} \) \(16.78 miles\)

OpenStudy (anonymous):

thank you

ganeshie8 (ganeshie8):

yw :)